# Terminal 3-folds that are not Cohen-Macaulay

### v1

@@ 5 The example over the 2-adic integers
The norm $N(x_{1})$ is a function on $U/G$ that defines a positive multiple of the divisor $F_{1}$, and so $F_{1}$ is Q-Cartier. Since the fixed point scheme $U^{G}$ is generically the Cartier divisor $E_{1}$, the ramification divisor of $f$ is $(p-1)E_{1}$, as mentioned above. That is, $K_{U}=f^{*}K_{U/G}+(p-1)E_{1}$. Also, the fixed point scheme is generically $E_{1}=\{x_{1}=0\}$ with coefficient 1, whereas $I(x_{1})$ vanishes to order 2 along $E_{1}$; so section 3 gives that the ramification of $f$ along $E_{1}$ is fierce. In particular, the ramification index $e$ is 1, meaning that $f^{*}F_{1}=E_{1}$. ∎
@@ 6 Characteristic 3
We work throughout over $k=\text{\bf F}_{3}$. Write $G=\text{\bf Z}/3=\langle\sigma:\sigma^{3}=1\rangle$, with $\tau:=\sigma^{-1}$. Let $Y_{0}=\text{\bf P}^{2}\times\text{\bf P}^{1}$ and $X=Y_{0}/G$. The only fixed point of $G$ on $Y_{0}$ is $A=([1,1,1],[0,1])$. So $X$ is normal of dimension 3, and $X$ is smooth over $k$ outside the image $P$ of $A$. Also, $3K_{X}$ is Cartier. By Fogarty, since $A$ is an isolated fixed point of $G=\text{\bf Z}/p$ on a smooth 3-fold in characteristic $p$, $X$ is not Cohen-Macaulay at $P$ [12, Proposition 4].
@@
First work over the chart $\{y_{1}=1\}$ in $Y_{1}$. We wrote out the $G$-action on this chart above, with coordinates $(y_{0},x_{1},y_{2})$. We defined $Y_{2}$ as the blow-up of $Y_{1}$ along the $G$-fixed curve $\{0=y_{0}=x_{1}\}$. So, over this open subset of $Y_{1}$, $Y_{2}$ is an open subset of
@@
To see all of $E_{1}$ in $Y_{2}$, we also need to work over $\{y_{2}=1\}\subset Y_{1}$, the chart containing the lower left vertex of $E_{1}$ in Figure 2. The corresponding open subset of $Y_{2}$ is an open subset of
@@
First consider the chart $\{z_{0}=1\}$ in $Y_{2}$. Then $(y_{0},y_{1},x_{2})=(y_{0},y_{1},y_{0}z_{2})$, and $G$ acts by
@@
The divisor class $\pi^{*}K_{Y_{0}/G}=K_{Y_{2}/G}-2F_{0}-F_{1}$ has some non-integer discrepancies, for example over the origin in the chart $\{z_{2}=1\}$ in $Y_{2}$ over $\{y_{2}=1\}$ in $Y_{1}$. As a result, $K_{Y_{0}/G}$ is not Cartier (as one can also check directly), in contrast to our examples in residue characteristic 2 ([26, Theorem 5.1] and Theorem 5.1). I expect that there is also a 3-fold $X$ over $\text{\bf F}_{3}$ that is terminal and non-Cohen-Macaulay with $K_{X}$ Cartier. Namely, one should replace $\text{\bf P}^{1}$ in Theorem 6.1 by the “Harbater-Katz-Gabber” curve $C$$\{0=y^{p}-yz^{p-1}-x^{p-1}z\}$ in $\text{\bf P}^{2}_{\text{\bf F}_{p}}$ [8], here with $p=3$, which has a $\text{\bf Z}/p$-action by $\tau([x,y,z])=[x+z,y,z]$ that preserves a nonzero 1-form near the fixed point $[1,0,0]$. For $p=3$, $C$ is a supersingular elliptic curve.
@@ 7 The example over the 3-adic integers
It remains to show that $X$ is terminal. One can resolve the singularities of $X$ by performing $G$-equivariant blow-ups of $Y_{0}$. However, as in sections 5 and 6, we will shorten the proof by recognizing that, after two $G$-equivariant blow-ups $Y_{2}\to Y_{1}\to Y_{0}$, the singularities of $Y_{2}/G$ become toric (in Kato’s mixed characteristic sense), namely quotients of a regular scheme by $\mu_{3}$. That makes it easy to check that $Y_{0}/G$ is terminal, without having to continue making $G$-equivariant blow-ups.
@@
The exceptional divisor $E_{0}$ is $\{x_{0}=0\}$ (isomorphic to $A^{2}_{\text{\bf F}_{3}}$, in this open set). Here $G$ acts by
@@
Finally, consider the chart $\{y_{2}=1\}$ in $Y_{1}$. Here $(x_{0},x_{1},e_{2})=(e_{2}y_{0},e_{2}y_{1},e_{2})$, and $G$ acts on the open set
@@ 8 Characteristic 5
We define the quintic del Pezzo surface (over any field) as the moduli space $\overline{M_{0,5}}$ of 5-pointed curves of genus 0. That makes it clear that the symmetric group $\Sigma_{5}$ acts on this surface.
@@
We work throughout over $k=\text{\bf F}_{5}$. Write $G=\text{\bf Z}/5=\langle\sigma:\sigma^{5}=1\rangle$, with $\tau:=\sigma^{-1}$. Let $Y_{0}=S_{5}\times\text{\bf P}^{1}$ and $X=Y_{0}/G$. In characteristic 5, $G$ has only one fixed point in $S_{5}$, and so $G$ has only one fixed point $P$ in $Y_{0}$. So $X$ is normal of dimension 3, and $X$ is smooth over $k$ outside the image of $P$ (which we also call $P$). Also, $5K_{X}$ is Cartier. By Fogarty, since $A$ is an isolated fixed point of $G=\text{\bf Z}/p$ on a smooth 3-fold in characteristic $p$, $X$ is not Cohen-Macaulay at $P$ [12, Proposition 4].
It remains to show that $X$ is terminal. This example is more complicated than those in characteristics 2 and 3, and it may be impossible to resolve the singularities of $X$ by performing $G$-equivariant blow-ups of $Y_{0}$. (Indeed, in the simpler situation of actions of $G=\text{\bf Z}/p$ in characteristic zero, one cannot always resolve the singularities of a quotient $Y/G$ via $G$-equivariant blow-ups of $Y$ when $p\geq 5$ [17, Claim 2.29.2].) Fortunately, as in earlier sections, we can reach toric singularities after some $G$-equivariant blow-ups. It will then be easy to check that $Y_{0}/G$ is terminal.
@@
It turns out that the fixed point set of $G$ on $Y_{1}$ is a curve isomorphic to $\text{\bf P}^{1}$. To check that, first work in the open subset $\{y_{0}=1\}$ in $Y_{1}$. Here $(s_{0},s_{1},s_{2})=(s_{0},s_{0}y_{1},s_{0}y_{2})$, and $G$ acts by
@@
The exceptional divisor $E_{0}$ is $\{s_{0}=0\}$, in this chart. The fixed point scheme $Y_{1}^{G}$ is defined by the vanishing of: $I(s_{0})=s_{0}(-y_{1}+O(s_{0}))$, $I(y_{1})=(y_{1}^{2}+O(s_{0}))/(1-y_{1}+O(s_{0}))$, and $I(y_{2})=y_{2}(y_{1}+O(s_{0}))/(1-y_{1}+O(s_{0}))$. We know that $Y_{1}^{G}$ is contained (as a set) in $E_{0}$ (since $Y_{0}^{G}$ is only the origin). So the fixed point set is the line $\{0=s_{0}=y_{1}\}$, in this chart.
In the chart $\{y_{1}=1\}$ in $Y_{1}$, we have $s_{0}=s_{1}y_{0}$ and $s_{2}=s_{1}y_{2}$, so we have coordinates $(y_{0},s_{1},y_{2})$. Here $E_{0}=\{s_{1}=0\}$. We can write the action of $G$ in these coordinates (for example using Magma). We find that the fixed point scheme $Y_{1}^{G}$ is defined by the vanishing of: $I(y_{0})=-1+O(s_{1})$, $I(s_{1})=s_{1}^{2}(1+y_{0}-2y_{0}^{2})$, and $I(y_{2})=s_{1}y_{2}(-1-y_{0}-y_{2}+2y_{0}^{2})$. Since $Y_{1}^{G}$ is contained (as a set) in $E_{0}$, the first equation shows that $Y_{1}^{G}$ is empty, in this chart. In the last chart $\{y_{2}=1\}$ in $Y_{1}$, we have coordinates $(y_{0},y_{1},s_{2})$, and $E_{0}=\{s_{2}=0\}$. The fixed point scheme is defined by: $I(y_{0})=-y_{1}+O(s_{2})$, $I(y_{1})=s_{2}(y_{1}+y_{1}^{2}+y_{0}y_{1}-2y_{0}^{2}+O(s_{2}))$, and $I(s_{2})=s_{2}^{2}(-1+O(s_{2}))$. Since $Y_{1}^{G}$ is contained (as a set) in $E_{0}$, the fixed point set is the line $\{0=y_{1}=s_{2}\}$, the same line seen in an earlier chart.
@@
To analyze the bad point above, change coordinates temporarily by $t_{1}=v_{1}+2$; then the bad point becomes the origin in coordinates $(w_{0},t_{1},y_{2})$. In these coordinates, we have $I(w_{0})=w_{0}^{2}(1-2t_{1}+O(w_{0}))$, $I(t_{1})=I(v_{1})=(-t_{1}-2t_{1}^{2}+O(w_{0}))$, and $I(y_{2})=w_{0}y_{2}(-2+O(w_{0}))$. Theorem 2.2 applies, with $s=e=w_{0}$, and we read off that $Y_{3}/G$ has singularity $\frac{1}{5}(1,-1,-2)$ at this point. That is terminal by the Reid-Tai criterion (Theorem 1.1, which will be convenient.
@@ 9 The example over the 5-adic integers
Let the group $G=\text{\bf Z}/5$ with generator $\tau$ act on the quintic del Pezzo surface $S_{5}$ over $\text{\bf Z}_{5}$ by an embedding of $G$ into the symmetric group $\Sigma_{5}=\operatorname{Aut}(S_{5})$. Let $R=\text{\bf Z}_{5}[e]/(e^{5}-5e^{4}+25e^{2}-25e+5)$, which is the ring of integers in a Galois extension of $\text{\bf Q}_{5}$ with group $G=\text{\bf Z}/5$. Let $G$ act on the scheme $(S_{5})_{R}$ by the diagonal action on $S_{5}$ and on $R$. Then the scheme $(S_{5})_{R}/G$ is terminal, not Cohen-Macaulay, of dimension 3, and flat over $\text{\bf Z}_{5}$.
We define the quintic del Pezzo surface $S_{5}$ (over any commutative ring) as the moduli space $\overline{M_{0,5}}$ of 5-pointed curves of genus 0. That makes it clear that the symmetric group $\Sigma_{5}$ acts on $S_{5}$.
@@
Over the open set $\{y_{0}=1\}$ in $Y_{1}$, $Y_{2}$ is the blow-up along the $G$-fixed curve $\{0=s_{0}=y_{1}\}$, so $Y_{2}$ has coordinates $((s_{0},y_{1},y_{2}),[w_{0},w_{1}])$. First take $\{w_{0}=1\}$, so $y_{1}=s_{0}w_{1}$, and we have coordinates $(s_{0},w_{1},y_{2})$. As in every other chart, there are three variables over $\text{\bf Z}_{5}$, but this is a regular scheme of dimension 3 because of the equation satisfied by $e$. In this case, we have $e_{2}=s_{0}y_{2}$, and so
@@
To analyze the bad point above, change coordinates temporarily by $t_{1}=v_{1}+2$; then the bad point becomes the origin in coordinates $(w_{0},t_{1},y_{2})$. In these coordinates, we have $I(w_{0})=w_{0}^{2}(1-2t_{1}+O(w_{0}))$, $I(t_{1})=I(v_{1})=(-t_{1}-2t_{1}^{2}+O(w_{0}))$, and $I(y_{2})=w_{0}y_{2}(-2+O(w_{0}))$. Theorem 2.2 applies, with $s=e=w_{0}$, and we read off that $Y_{3}/G$ has singularity $\frac{1}{5}(1,-1,-2)$ at this point. That is terminal by the Reid-Tai criterion (Theorem 1.1, which will be convenient.

### v2

@@ 5 The example over the 2-adic integers
The norm $N(x_{1})$ is a function on $U/G$ that defines a positive multiple of the divisor $F_{1}$, and so $F_{1}$ is Q-Cartier. Since the fixed point scheme $U^{G}$ is generically the Cartier divisor $E_{1}$, the ramification divisor of $f$ is $(p-1)E_{1}$ by Lemma 3.1. That is, $K_{U}=f^{*}K_{U/G}+(p-1)E_{1}$. Also, the fixed point scheme is generically $E_{1}=\{x_{1}=0\}$ with coefficient 1, whereas $I(x_{1})$ vanishes to order 2 along $E_{1}$; so section 3 gives that the ramification of $f$ along $E_{1}$ is fierce. In particular, the ramification index $e$ is 1, meaning that $f^{*}F_{1}=E_{1}$. ∎
@@ 6 Characteristic 3
We work throughout over $k=\text{\bf F}_{3}$. Write $G=\text{\bf Z}/3=\langle\sigma:\sigma^{3}=1\rangle$, with $\tau:=\sigma^{-1}$. Let $Y_{0}=\text{\bf P}^{2}\times\text{\bf P}^{1}$ and $X=Y_{0}/G$. The only fixed point of $G$ on $Y_{0}$ is $P=([1,1,1],[0,1])$. So $X$ is normal of dimension 3, and $X$ is smooth over $k$ outside the image of $P$, which we also call $P$. Also, $3K_{X}$ is Cartier. By Fogarty, since $P$ is an isolated fixed point of $G=\text{\bf Z}/p$ on a smooth 3-fold in characteristic $p$, $X$ is not Cohen-Macaulay at $P$ [13, Proposition 4].
@@
First work over the chart $\{y_{1}=1\}$ in $Y_{1}$. We wrote out the $G$-action on this chart above, with coordinates $(y_{0},x_{1},y_{2})$. Here $E_{0}=\{x_{0}=0\}$. We defined $Y_{2}$ as the blow-up of $Y_{1}$ along the $G$-fixed curve $\{0=y_{0}=x_{1}\}$ in $E_{0}$. So, over this open subset of $Y_{1}$, $Y_{2}$ is an open subset of
@@
To see all of $E_{1}$ in $Y_{2}$, we also need to work over $\{y_{2}=1\}\subset Y_{1}$. The coordinates are $(y_{0},y_{1},x_{2})$, and $E_{0}=\{x_{2}=0\}$. The corresponding open subset of $Y_{2}$ is an open subset of
@@
First consider $\{z_{0}=1\}$ in $Y_{2}$, the chart containing the lower left vertex of $E_{1}$ in Figure 2. Then $(y_{0},y_{1},x_{2})=(y_{0},y_{1},y_{0}z_{2})$, and $G$ acts by
@@
The divisor class $\pi^{*}K_{Y_{0}/G}=K_{Y_{2}/G}-2F_{0}-F_{1}$ has some non-integer discrepancies, for example over the origin in the chart $\{z_{2}=1\}$ in $Y_{2}$ over $\{y_{2}=1\}$ in $Y_{1}$. As a result, $K_{Y_{0}/G}$ is not Cartier (as one can also check directly), in contrast to our examples in residue characteristic 2 ([29, Theorem 5.1] and Theorem 5.1). I expect that there is also a 3-fold $X$ over $\text{\bf F}_{3}$ that is terminal and non-Cohen-Macaulay with $K_{X}$ Cartier. Namely, one should replace $\text{\bf P}^{1}$ in Theorem 6.1 by the “Harbater-Katz-Gabber” curve $C=\{0=y^{p}-yz^{p-1}-x^{p-1}z\}$ in $\text{\bf P}^{2}_{\text{\bf F}_{p}}$ [9], here with $p=3$, which has a $\text{\bf Z}/p$-action by $\tau([x,y,z])=[x+z,y,z]$ that preserves a nonzero 1-form near the fixed point $[1,0,0]$. For $p=3$, $C$ is a supersingular elliptic curve.
@@ 7 The example over the 3-adic integers
It remains to show that $X$ is terminal. One can resolve the singularities of $X$ by performing $G$-equivariant blow-ups of $Y_{0}$. However, as in sections 5 and 6, we will shorten the proof by recognizing that, after two $G$-equivariant blow-ups $Y_{2}\to Y_{1}\to Y_{0}$, the singularities of $Y_{2}/G$ become toric (in Kato’s mixed-characteristic sense), namely quotients of a regular scheme by $\mu_{3}$. That makes it easy to check that $Y_{0}/G$ is terminal, without having to continue making $G$-equivariant blow-ups.
@@
The exceptional divisor $E_{0}$ is $\{x_{0}=0\}$. Here $G$ acts by
@@
Finally, consider the chart $\{y_{2}=1\}$ in $Y_{1}$. We have coordinates $(y_{0},y_{1},e_{2})$, with $(x_{0},x_{1},e_{2})=(e_{2}y_{0},e_{2}y_{1},e_{2})$, and $E_{0}=\{e_{2}=0\}$. Here $G$ acts on the open set
@@ 8 Characteristic 5
We define the quintic del Pezzo surface (over any field) as the moduli space $\overline{M_{0,5}}$ of 5-pointed stable curves of genus 0. That makes it clear that the symmetric group $\Sigma_{5}$ acts on this surface.
@@
We work throughout over $k=\text{\bf F}_{5}$. Write $G=\text{\bf Z}/5=\langle\sigma:\sigma^{5}=1\rangle$, with $\tau:=\sigma^{-1}$. Let $Y_{0}=S_{5}\times\text{\bf P}^{1}$ and $X=Y_{0}/G$. In characteristic 5, $G$ has only one fixed point in $S_{5}$, and so $G$ has only one fixed point $P$ in $Y_{0}$. So $X$ is normal of dimension 3, and $X$ is smooth over $k$ outside the image of $P$ (which we also call $P$). Also, $5K_{X}$ is Cartier. By Fogarty, since $P$ is an isolated fixed point of $G=\text{\bf Z}/p$ on a smooth 3-fold in characteristic $p$, $X$ is not Cohen-Macaulay at $P$ [13, Proposition 4].
It remains to show that $X$ is terminal. This example is more complicated than those in characteristics 2 and 3, and it may be impossible to resolve the singularities of $X$ by performing $G$-equivariant blow-ups of $Y_{0}$. (Indeed, in the simpler situation of actions of $G=\text{\bf Z}/p$ in characteristic zero, one cannot always resolve the singularities of a quotient $Y/G$ via $G$-equivariant blow-ups of $Y$ when $p\geq 5$ [19, Claim 2.29.2].) Fortunately, as in earlier sections, we can reach toric singularities after some $G$-equivariant blow-ups. It will then be easy to check that $Y_{0}/G$ is terminal.
@@
The exceptional divisor $E_{0}$ is isomorphic to $\text{\bf P}^{2}$. It turns out that the fixed point set of $G$ on $Y_{1}$ is a curve isomorphic to $\text{\bf P}^{1}$ in $E_{0}$. To check that, first work in the open subset $\{y_{0}=1\}$ in $Y_{1}$. Here $(s_{0},s_{1},s_{2})=(s_{0},s_{0}y_{1},s_{0}y_{2})$, and $G$ acts by
@@
Here $E_{0}=\{s_{0}=0\}$. The fixed point scheme $Y_{1}^{G}$ is defined by the vanishing of: $I(s_{0})=s_{0}(-y_{1}+O(s_{0}))$, $I(y_{1})=(y_{1}^{2}+O(s_{0}))/(1-y_{1}+O(s_{0}))$, and $I(y_{2})=y_{2}(y_{1}+O(s_{0}))/(1-y_{1}+O(s_{0}))$. We know that $Y_{1}^{G}$ is contained (as a set) in $E_{0}$ (since $Y_{0}^{G}$ is only the origin). So the fixed point set is the line $\{0=s_{0}=y_{1}\}$, in this chart.
In the chart $\{y_{1}=1\}$ in $Y_{1}$, we have $s_{0}=s_{1}y_{0}$ and $s_{2}=s_{1}y_{2}$, so we have coordinates $(y_{0},s_{1},y_{2})$. Here $E_{0}=\{s_{1}=0\}$. We can write the action of $G$ in these coordinates (for example using Magma). We find that the fixed point scheme $Y_{1}^{G}$ is defined by: $I(y_{0})=-1+O(s_{1})$, $I(s_{1})=s_{1}^{2}(1+y_{0}-2y_{0}^{2})$, and $I(y_{2})=s_{1}y_{2}(-1-y_{0}-y_{2}+2y_{0}^{2})$. Since $Y_{1}^{G}$ is contained (as a set) in $E_{0}$, the first equation shows that $Y_{1}^{G}$ is empty, in this chart. In the last chart $\{y_{2}=1\}$ in $Y_{1}$, we have coordinates $(y_{0},y_{1},s_{2})$, and $E_{0}=\{s_{2}=0\}$. The fixed point scheme is defined by: $I(y_{0})=-y_{1}+O(s_{2})$, $I(y_{1})=s_{2}(y_{1}+y_{1}^{2}+y_{0}y_{1}-2y_{0}^{2}+O(s_{2}))$, and $I(s_{2})=s_{2}^{2}(-1+O(s_{2}))$. Since $Y_{1}^{G}$ is contained (as a set) in $E_{0}$, the fixed point set is the line $\{0=y_{1}=s_{2}\}$, the same line seen in an earlier chart.
@@
To analyze the bad point above, change coordinates temporarily by $t_{1}=v_{1}+2$; then the bad point becomes the origin in coordinates $(w_{0},t_{1},y_{2})$. In these coordinates, we have $I(w_{0})=w_{0}^{2}(1-2t_{1}+O(w_{0}))$, $I(t_{1})=I(v_{1})=(-t_{1}-2t_{1}^{2}+O(w_{0}))$, and $I(y_{2})=w_{0}y_{2}(-2+O(w_{0}))$. Theorem 2.2 applies, with $s=e=w_{0}$, and we read off that $Y_{3}/G$ has singularity $\frac{1}{5}(1,-1,-2)$ at this point. That is terminal, by the Reid-Tai criterion (Theorem 1.1).
@@ 9 The example over the 5-adic integers
Let the group $G=\text{\bf Z}/5$ act on the quintic del Pezzo surface $S_{5}$ over $\text{\bf Z}_{5}$ by an embedding of $G$ into the symmetric group $\Sigma_{5}=\operatorname{Aut}(S_{5})$. Let $R=\text{\bf Z}_{5}[e]/(e^{5}-5e^{4}+25e^{2}-25e+5)$, which is the ring of integers in a Galois extension of $\text{\bf Q}_{5}$ with group $G=\text{\bf Z}/5$. Let $G$ act on the scheme $(S_{5})_{R}$ by the diagonal action on $S_{5}$ and on $R$. Then the scheme $(S_{5})_{R}/G$ is terminal, not Cohen-Macaulay, of dimension 3, and flat over $\text{\bf Z}_{5}$.
We define the quintic del Pezzo surface $S_{5}$ (over any commutative ring) as the moduli space $\overline{M_{0,5}}$ of 5-pointed stable curves of genus 0. That makes it clear that the symmetric group $\Sigma_{5}$ acts on $S_{5}$.
@@
Over the open set $\{y_{0}=1\}$ in $Y_{1}$, $Y_{2}$ is the blow-up along the $G$-fixed curve $\{0=s_{0}=y_{1}\}$, so $Y_{2}$ has coordinates $((s_{0},y_{1},y_{2}),[w_{0},w_{1}])$. First take $\{w_{0}=1\}$, so $y_{1}=s_{0}w_{1}$, and we have coordinates $(s_{0},w_{1},y_{2})$. As in every other chart, there are three variables over $\text{\bf Z}_{5}$, but this is a regular scheme of dimension 3 because of the equation satisfied by $e_{2}$. In this case, we have $e_{2}=s_{0}y_{2}$, and so
@@
To analyze the bad point above, change coordinates temporarily by $t_{1}=v_{1}+2$; then the bad point becomes the origin in coordinates $(w_{0},t_{1},y_{2})$. In these coordinates, we have $I(w_{0})=w_{0}^{2}(1-2t_{1}+O(w_{0}))$, $I(t_{1})=I(v_{1})=(-t_{1}-2t_{1}^{2}+O(w_{0}))$, and $I(y_{2})=w_{0}y_{2}(-2+O(w_{0}))$. Theorem 2.2 applies, with $s=e=w_{0}$, and we read off that $Y_{3}/G$ has singularity $\frac{1}{5}(1,-1,-2)$ at this point. That is terminal, by the Reid-Tai criterion (Theorem 1.1).